can you help me convert to slope intercept;
y-2=-1/5(x-4)

Found 2 solutions by fractalier, Alan3354:
Answer by fractalier(6550)

From
y-2=-1/5(x-4)
let us distribute and get
y - 2 = (-1/5)x + 4/5
Now add two to both sides (2 = 10/5)
y = (-1/5)x + 14/5

Answer by Alan3354(69443)

can you help me convert to slope intercept;
y-2=-1/5(x-4)
==============
Solve for y.